A ring that has radius a lies in the z 0 plane with its center at the origin. The ring is uniformly charged and has a total charge Q (a) Find Ez on the z axis at 0.1a. 2 (b) Find Ez on the z axis at 0.4a. kQ (c) Find Ez on the z axis at 0.7a. ko (d) Find Ez on the z axis at a. ko (e) Find Ez on the z axis at 1.6a kQ. 2 (f) Use your results to plot Ez versus z for both positive and negative values of z. (Assume that these distances are exact. Do this on paper. Your instructor may ask you to turn in this work.) eBook
Answer #1

answer) we know the formula

Ez=kQz/z2+a2)3/2

a) Ez(0.1,a)=kQ*0.1a/((0.1a)2+(a2))3/2=0.099kQ/a2

b)Ez(0.4a,a)=kQ0.4a/((0.4a2+a2)3/2=0.377kQ/a2

c)Ez(0.7,a)=kQ0.7a/((0.7a)2+a2)3/2=0.385kQ/a2

d) Ez(a)=kQa/(a2+a2)3/2=0.354kQ/a2

e)Ez(1.6a)=kQ1.6a/(1.6a)2+a2)3/2=0.238kQ/a2

f) the figure is below2 2 43210 1.234- O 凸